at maximum speed, an airplane travels 2400 miles against the wind in 6 hours. Flying with the wind, the plane can travel the same distance in 5 hours.
what is the speed of the airplane with no wind?

Answer by mananth(16949)

let speed of plane in still air bex
and wind rate be y
2400 / x-y = 6
2400/6 = x-y
400 = x-y
2400 / x+y=5
x+y = 480
x-y+ x+y = 480+400
2x=880
x=440 mph rate of plane in still air
so y= 40 mph rate of wind